OpenStudy (anonymous):

Which answer shows this equation in standard form? 7 – 3(x – y) = –5x + 2 A. –8x + 3y = –5 B. –2x – 3y = 5 C. 2x + 3y = –5 D. 2x – 3y = –5

OpenStudy (anonymous):

the thing you wrote is equal to 7 – 3x – 3y = –5x + 2 this is the same as 7 – 3y = –2x + 2 this is the same as 5=-2x+3y

OpenStudy (anonymous):

which is the same as 2x-3y=-5

OpenStudy (anonymous):

which is D

OpenStudy (anonymous):

thx so much